OpenStudy (anonymous):

A phone company charges a connection fee of $0.92 and a variable cost per minute of $0.06 for a call. Assume the number of minutes is your independent variable (x) and the cost is your dependent variable (y). Find 'y', the cost of a phone call that lasts x minutes. Write your equation in the form y = _____________.

OpenStudy (akashdeepdeb):

Answer this: If I talk for 10 minutes with my friend. What will be the charge? Well, I'd have a mandatory charge = 0.92 Charge per minute = 0.06 per minute for 10 minutes = 0.06 * 10 minutes = 0.6$ Thus total charge for 10 minutes would be = 0.92 + 0.6 = 1.52$ Now what if I talked for 11 minutes? 12 minutes? 100 minutes? 1 hour? x minutes? Thus we take the unknown number of minutes as x. Total cost = Mandatory cost + Cost per minute * x = y = 0.92 + 0.06x Getting this? :)